OpenStudy (anonymous):

write in standard form an equation of the line that passes through the two points. use integer coefficients. FOR (4,9) and (-2,-6)

OpenStudy (anonymous):

find the slope first

OpenStudy (anonymous):

to find the slope you use (y2-y1)/(x2-x1) you'll get 15/6 put that into the equation y=mx+b and use one of the ordered pairs to put in a y and x value t find what b is

OpenStudy (l):

(-6-9)/(-2-4)=(y-9)/(x-4) 5x-2y-2=0

OpenStudy (anonymous):

THANK YOU

OpenStudy (anonymous):

(-6-9)/(-2-4) = -15/-6 = 15/6 y=mx+b y=(15/6)x+b to find b i'll use the coordinates (4,9) 9=(15/6)(4)+b 9=10+b subtracting 10 from both sides -1=b so mow you have the equation y=(15/6)x - 1 to turn this into standard form (-15/6)x + y = -1 divide by -6 15x-6y=6 is that what you got?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay, if you have an answer key (like at the back of your textbook) check it to make sure you have the right thing

OpenStudy (anonymous):

can you help me with another one please?

OpenStudy (anonymous):

post it and we'll see

OpenStudy (anonymous):

write in standard form an equation of the line that passes through the given point and has the given slope. use integer coefficients. FOR (2,3) m = -4

OpenStudy (anonymous):

okay, so put this into the equation y=mx+b again to find b you'll get 11 and have the equation y=-4x+11, in standard form this is 4x+y=11 is that what you have?

OpenStudy (anonymous):

no i did something wrong

OpenStudy (anonymous):

\[y=mx+b\] substitute the x and y for their values (the coordinates) 3=-4(2)+b 3=-8+b add 8 to both sides 11=b now you have, in slope intercept form, y=-4x+11 in standard form, this is 4x+y=11